### v2

@@ 4.4.2 Heterogeneous Multi-modal
Heterogeneous multi-modal paradigm refers to the tasks where the input data consists of multiple modalities with different data types. In the field of text-driven generation, Mamba architectures are utilized for visual processing. To enhance the generalization capabilities of image editing approaches, InstructGIE (Meng et al, 2024) reformed the conditional latent diffusion model by integrating a vision encoder based on VMamba, and used the VMamba layer to inject the processed visual condition information into the frozen Stable Diffusion (Rombach et al, 2022) model. In the realm of Multimodal large language models (MLLMs), VLMamba (Qiao et al, 2024) replaced the transformer-based backbone language model such as LLama (Touvron et al, 2023) with the pre-trained Mamba language model. Additionally, it introduced the Multi-Modal Connector, comprising a Vision Selective Scan (VSS) module and two linear layers. Meteor (Lee et al, 2024) introduced multifaceted rationales and applied the Mamba architecture to embed these lengthy rationales to enhance the understanding and answering capabilities of MLLMs. Cobra (Zhao et al, 2024a) introduced a linear computational complexity MLLM, utilizing DINOv2 (Oquab et al, 2023) and SigLIP (Zhai et al, 2023) as vision encoder and directly employing the Mamba language model as the backbone. ReMamber (Yang et al, 2024d) is a novel referring image segmentation framework containing several Mamba Twister blocks. Each block comprises Visual State Space (VSS) layers to extract the vision feature and a twister layer to integrate the textual information into the visual modality. CLIP-Mamba (Huang et al, 2024f) is the first to train a transferable Mamba model using Contrastive Language-Image Pre-training (CLIP). It demonstrated that compared to ViT, Mamba is more parameter-efficient in zero-shot classification tasks and exhibits exceptional robustness in out-of-distribution (OOD) tasks. Additionally, Hessian visualizations suggest that Mamba models may face challenges in optimization. SurvMamba (Chen et al, 2024g) was designed to enhance the accuracy of survival prediction by integrating pathological images and genomic data. It introduced the Hierarchical Interaction Mamba (HIM) to facilitate efficient intra-modal interactions at different granularities, and the Interaction Fusion Mamba to aggregate the cascaded multi-modal features. RSCaMa (Liu et al, 2024a) employed multiple CaMa layers between the backbone and the language decoder to facilitate efficient joint spatial-temporal modeling for remote sensing image change captioning.
SpikeMba (Li et al, 2024e) combined the Mamba and Spiking mechanism for temporal video grounding. It comprises the Mamba-based Contextual Moment Reasoner with learnable relevant slots, Spiking Saliency Detector, and Multi-modal Relevant Mamba. BroadMamba (Shou et al, 2024) applied the Mamba architecture to multimodal emotion recognition in conversation, and designed a Broad Mamba using bidirectional SSM for the feature disentanglement stage, enabling the extraction of long-range contextual semantic information from different modalities. Coupled Mamba (Li et al, 2024f) proposed coupling state chains of inter-modalities while maintaining the independence of intra-modality state processes. It also derived the global convolution kernel to enable parallelism for the coupled state transition scheme.
Mamba-FETrack (Huang et al, 2024a) introduced a RGB-Event tracking framework that integrates video frames and event streams for target object localization. The framework employed the Mamba architecture for modality-specific feature extraction and multi-modal feature fusion. Motion Mamba (Zhang et al, 2024e) incorporated two modules based on the vanilla Mamba into a diffusion-based generative system for text-to-motion synthesis, namely hierarchical temporal Mamba and bidirectional spatial Mamba. TM-Mamba (Wang et al, 2024f) introduced a novel text-based motion grounding task and employed the bidirectional SSM to construct the text-controlled selection mechanism, dynamically incorporating global temporal information from the motion sequence based on the text query.
5 Challenges and Future Directions
@@
5.1 Scalability
5.1.1 Model Size
@@
Firstly, the Mamba architecture faces stability issues when scaled to large network configurations (Patro and Agneeswaran, 2024a). The underlying causes of Mamba’s instability remain inadequately understood. This instability often manifests as vanishing or exploding gradients (Patro and Agneeswaran, 2024b), leading to performance degradation or even training collapse. To address the stability issues, some studies utilize stable reparameterization to balance gradient scales in state space models (Wang and Li, 2024). Some works utilize learnable Fourier Transforms to adjust the eigenvalues of the evolution matrix, ensuring they are negative real numbers, which helps maintain stability (Patro and Agneeswaran, 2024b). Despite these efforts, most visual Mamba models are limited to base or smaller scales, which may restrict their performance.
@@
Secondly, visual Mamba models, when controlled for similar parameters or FLOPs, are generally outperformed by cutting-edge CNNs, advanced visual Transformers, and hybrid architectures that integrate multiple model types. This performance gap also hinders the scaling of visual Mamba models and underscores the necessity to enhance their performance. Integrating architectures or techniques proven effective in CNNs and visual Transformers could help bridge the performance gap. Besides, while there may be some compromise in performance, the non-hierarchical architectures comprising identical blocks facilitate optimization for hardware acceleration. This aspect is crucial for scaling to large models and enhancing computational efficiency. In addition, self-supervised pre-training plays a critical role in enhancing the model performance. For instance, ARM (Ren et al, 2024b) demonstrates that the autoregressive pre-training can significantly improve the performance of the Mamba architecture and address the instability issues, efficiently unlocking its potential for scaling to very large model sizes. However, the self-supervised pre-training strategies for Mamba models are still nascent.
@@
Future research should focus on optimizing these aspects to harness the full capabilities of large network configurations.
5.1.2 Data
@@
The Mamba architecture theoretically reduces computational complexity, which enhances its efficiency for processing long sequence data. Moreover, unlike many sequence models that do not benefit from extended contexts (Shi et al, 2023), Mamba can selectively disregard irrelevant information, resulting in consistent performance improvements as context length increases. The efficiency and efficacy of Mamba architecture bring significant opportunities for data scalability.
@@
Furthermore, given the parameter efficiency of the Mamba models and their inherent local inductive bias, which typically mitigate overfitting, Mamba models have considerable potential for achieving optimal performance without relying on large-scale datasets. This opens up opportunities for developing small yet effective Mamba-based models advantageous in resource-constrained environments, a pathway that has not been extensively explored in current research.
@@ 5.1.3 Hardware-aware
Mamba uses kernel fusion and recomputation to make the computation of the parallel SSM scan fast and memory-efficient. Specisfically, the discretization step, the scan, and the multiplication with the projection matrix are fused into one kernel. Then the SSM parameters are loaded from slow HBM to fast SRAM for calculating the fused kernel in SRAM. Subsequently, the outputs are written back to HBM. In addition, Mamba chooses to recompute the intermediate states for the backpropagation since the recomputation is faster than storing them and reading them from HBM. These two strategies accelerate the scan operation and reduce the GPU requirement.
However, according to research findings (Wang et al, 2024g; Dao and Gu, 2024), GPU consumption for the Mamba model does not consistently demonstrate a reduction compared to the Transformer model, highlighting the need for optimized, hardware-aware Mamba algorithms tailored for vision tasks. Mamba-2 (Dao and Gu, 2024) establishes theoretical connections between structured SSMs and variants of attention by reformulating different approaches to computing SSMs as various matrix multiplication algorithms on structured matrices and generalizing the linear attention (Katharopoulos et al, 2020) to structured masked attention. Based on this theoretical analysis, Mamba-2 introduces an efficient State Space Duality (SSD) algorithm for SSM computation and modifies the Mamba block to be tensor parallelism friendly. This framework allows Mamba-2 to leverage system optimizations developed for Transformers, including tensor parallelism and sequence parallelism, facilitating the parallel training of large models using multiple GPUs. Moreover, unlike Transformers, Mamba-2 can be efficiently trained with variable sequence lengths. However, the adaptation of Mamba-2 to vision tasks remains an area that requires further investigation.
@@ 5.1.4 Task
5.2 Causality
In visual Transformer-based architectures, the attention matrix explicitly captures dependencies between each pair of visual tokens, enabling predictions at any location to have a global receptive field. However, in Mamba-based architectures, the S6 blocks make predictions for each token based solely on the hidden state $h_{t-1}$ and the current input $x_{t}$. The $h_{t-1}$ only incorporates information from previously scanned tokens and disregards information from other tokens. Consequently, the vanilla Mamba is a causal system where predictions rely solely on the current and previous inputs. Although the causality is naturally suitable for autoregressive models in sequential vision tasks (Ren et al, 2024b), it contradicts the nature of spatial images, resulting in a fundamental mismatch between Mamba-based architectures and common visual tasks. To address this issue, existing works have proposed to integrate multiple scanning directions to ensure a global receptive field. While these strategies partially mitigate the undesirable causality of Mamba in vision tasks, they still encounter challenges or introduce new drawbacks that require resolution.
Firstly, while the utilization of multiple scanning directions helps establish a global receptive field, this approach does not effectively address the underlying causality issue. Tokens at different locations are still treated inconsistently due to their varying positions within the sequences, inevitably leading to a prioritization of more recently scanned tokens. Consequently, this strategy struggles to preserve spatial relationships and inadvertently introduces undesired bias for vision recognition as a result of the scanning order. In contrast, the self-attention mechanism in Transformers consistently consider all tokens to capture dependencies, showcasing its superiority to Mamba.
Secondly, the incorporation of multiple scanning strategies leads to redundancy and sub-optimal performance. To enhance Mamba’s perception at different locations, numerous approaches incorporate not only different scanning directions but also different scanning modes, axes, and continuity. While incorporating more scanning strategies may increase the information captured in specific locations, it introduces a significant amount of redundant information within these scanning sequences. This redundancy can impact model efficiency and hinder effective knowledge extraction for vision tasks.
@@ 5.3 In-context Learning
Currently, Transformer models are the primary large models possessing ICL capabilities. The efficiency and long sequence modeling ability of Mamba make it a promising alternative to Transformers for ICL. Furthermore, the inherent causality of the Mamba model eliminates the need for position embeddings, which are typically a constraint in extending the ICL capabilities of Transformers. In (Park et al, 2024), the authors investigate the ICL capabilities of Mamba compared to Transformers, focusing on regression ICL tasks and small-scale models. The results demonstrate that Mamba is capable of ICL and performs comparably to Transformers in standard regression ICL tasks. Additionally, Mamba outperforms Transformers in tasks such as sparse parity learning. However, Mamba struggles with tasks involving non-standard retrieval functionality, which Transformers can handle with ease. Research detailed in (Grazzi et al, 2024) further demonstrates Mamba’s capabilities on more complex NLP in-context tasks and shows that Mamba’s ICL efficacy scales positively with the number of in-context examples.
@@
5.4 Trustworthiness
5.4.1 Interpretability
Some studies have provided experimental evidence to elucidate the mechanisms underlying the Mamba model in NLP, focusing on its in-context learning capabilities (Grazzi et al, 2024; Park et al, 2024) and factual recall capabilities (Sharma et al, 2024). Additionally, other works (Cirone et al, 2024) have laid theoretical foundations for Mamba’s applications in NLP. Despite these advancements, explaining why Mamba performs effectively on visual tasks remains challenging. In (Ali et al, 2024), the authors reinterpret Mamba layers as self-attention mechanisms, thereby shedding light on its functionality both empirically and theoretically. However, the distinct learning characteristics of visual Mamba and its parallels with other foundational architectures, such as RNNs, CNNs, and ViTs, still demand deeper interpretation.
5.4.2 Generalization
According to findings in (Long et al, 2024), the hidden states in Mamba are likely to accumulate or even amplify domain-specific information, which can adversely affect its generalization performance. Moreover, the 1D scanning strategies inherent to the model may inadvertently capture domain-specific biases, and current scanning techniques often fail to address the need for domain-agnostic information processing. Addressing these challenges to better harness Mamba’s strengths while improving its generalization remains a critical area for future research.
5.4.3 Safety
Research presented in (Du et al, 2024) demonstrates VMamba’s (Liu et al, 2024g) strengths in adversarial resilience and general robustness. However, it also identifies limitations in scalability when dealing with these tasks. The study includes white-box attacks on VMamba to examine the behavior of its novel components under adversarial conditions. The findings indicate that while parameter $\Delta$ exhibits robustness, parameters $\mathbf{B}$ and $\mathbf{C}$ are susceptible to attacks. This differential vulnerability among parameters contributes to VMamba’s scalability challenges in maintaining robustness. Furthermore, the results reveal that VMamba exhibits particular sensitivity to interruptions in the continuity of its scanning trajectory and the integrity of spatial information. Enhancing the safety of visual Mamba remains an unresolved challenge in the field.
6 Conclusion

### v3

@@ 4.4.2. Heterogeneous Multi-modal
In the field of text-driven generation, Mamba architectures are utilized for visual processing. To enhance the generalization capabilities of image editing approaches, InstructGIE (Meng et al., 2024) reformed the conditional latent diffusion model by integrating a vision encoder based on VMamba (Liu et al., 2024e), and used the VMamba layer to inject the processed visual condition information into the frozen Stable Diffusion (Rombach et al., 2022) model. In the realm of MLLMs, Meteor (Lee et al., 2024) introduced multifaceted rationales and applied the Mamba architecture to embed these lengthy rationales to enhance the understanding and answering capabilities of MLLMs. RoboMamba (Liu et al., 2024d) aligned the vision encoder with an efficient Mamba language model to enable cost-effective robot manipulation capabilities. ReMamber (Yang et al., 2024c) introduced a novel referring image segmentation framework containing multiple Mamba twister blocks. Each block comprises VMamba (Liu et al., 2024e) layers to extract the vision feature and a twister layer to integrate the textual information into the visual modality. The Twister layer first constructs a hybrid feature cube and subsequently performs channel-wise and spatial scanning for feature fusion. RSCaMa (Liu et al., 2024a) employed multiple CaMa layers between the backbone and the language decoder to facilitate efficient joint spatial-temporal modeling for remote sensing image change captioning. Specifically, the CaMa layer enhances change awareness by multiplying the differencing features with the output of SSMs. It further employs a Temporal-first scanning strategy to facilitate effective temporal modeling.
CoupledMamba (Li et al., 2024k) proposed coupling state chains of inter-modalities while maintaining the independence of intra-modality state processes. It also derived the global convolution kernel to enable parallelism for the coupled state transition scheme. MotionMamba (Zhang et al., 2024f) incorporated two modules based on the vanilla Mamba into a diffusion-based generative system for text-to-motion synthesis, namely hierarchical temporal Mamba and bidirectional spatial Mamba. MambaGesture (Fu et al., 2024b) integrated Mamba architecture with the attention mechanism to form the basic block for the diffusion model in the co-speech gesture generation task.
5. Challenges and Future Directions
@@
5.1. Scalability
5.1.1. Model Size
@@
Firstly, the Mamba architecture faces stability issues when scaled to large network configurations (Patro and Agneeswaran, 2024a). The underlying causes of Mamba’s instability remain inadequately understood. This instability often manifests as vanishing or exploding gradients (Patro and Agneeswaran, 2024b), leading to performance degradation or even training collapse. To address the stability issues, some studies utilize stable reparameterization to balance gradient scales in state space models (Wang and Li, 2024; Ran-Milo et al., 2024; Orvieto et al., 2024). Some works utilize learnable Fourier Transforms to adjust the eigenvalues of the evolution matrix, ensuring they are negative real numbers, which helps maintain stability (Patro and Agneeswaran, 2024b). Mamba-2 incorporates an additional normalization layer in the block before the final output projection, mitigating instability issues encountered in large models (Dao and Gu, 2024a). Despite these efforts, most visual Mamba models are limited to base or smaller scales, which may restrict their performance.
@@
Secondly, visual Mamba models, when controlled for similar parameters or FLOPs, are generally outperformed by cutting-edge CNNs, advanced visual Transformers, and hybrid architectures that integrate multiple model types. This performance gap also hinders the scaling of visual Mamba models and underscores the necessity to enhance their performance. Integrating architectures or techniques proven effective in CNNs and visual Transformers could help bridge the performance gap. Besides, while there may be some compromise in performance, the non-hierarchical architectures comprising identical blocks facilitate optimization for hardware acceleration. This aspect is crucial for scaling to large models and enhancing computational efficiency. In addition, self-supervised pre-training plays a critical role in enhancing the model performance. For instance, ARM (Ren et al., 2024a) demonstrated that the autoregressive pre-training can significantly improve the performance of the Mamba architecture and address the instability issues, efficiently unlocking its potential for scaling to very large model sizes. However, the self-supervised pre-training strategies for Mamba models are still nascent.
More recently, Mamba-2 (Dao and Gu, 2024a) outlined several potential directions for further enhancing its performance. Given that SSD enhances hardware efficiency by limiting expressivity through the use of restricted structured matrices, future work could explore refining structured matrix algorithms toward more generalized diagonal SSMs to balance efficiency and expressivity (Merrill et al., 2024). Mamba-2 further suggests that advancements from a matrix mixer perspective could broaden the applications of sequence models, including the development of principled non-causal Mamba variants and methods to bridge the gap between softmax attention and sub-quadratic models by analyzing their matrix transformation structures.
@@
Future research should explore optimizing these aspects to harness the full capabilities of large network configurations.
5.1.2. Data
@@
The Mamba architecture theoretically reduces computational complexity, enhancing its efficiency in long sequence processing. Moreover, unlike many sequence models that do not benefit from extended contexts (Shi et al., 2023), Mamba can selectively disregard irrelevant information, yielding consistent performance improvements as context length increases. The efficiency and efficacy of Mamba architecture bring significant opportunities for data scalability.
@@
Furthermore, given the parameter efficiency of the Mamba models and their inherent local inductive bias, which typically mitigate overfitting, Mamba models have considerable potential for achieving optimal performance without relying on large-scale datasets. This opens up opportunities for developing small yet effective Mamba-based models advantageous in resource-constrained environments, a pathway that is underexplored in current research.
@@ 5.1.3. Hardware-aware
Mamba-1 uses kernel fusion and recomputation to make the computation of the parallel SSM scan fast and memory-efficient. Specifically, the discretization step, the scan, and the multiplication with the projection matrix are fused into one kernel. Then the SSM parameters are loaded from slow HBM to fast SRAM for calculating the fused kernel in SRAM. Subsequently, the outputs are written back to HBM. In addition, Mamba-1 chooses to recompute the intermediate states for the backpropagation since the recomputation is faster than storing them and reading them from HBM. These two strategies accelerate the scan operation and reduce the GPU requirement. However, according to research findings (Wang et al., 2024h; Dao and Gu, 2024a), GPU consumption for the Mamba model does not consistently demonstrate a reduction compared to the Transformer model, highlighting the need for optimized, hardware-aware Mamba algorithms tailored for vision tasks.
Mamba-2 (Dao and Gu, 2024a) establishes theoretical connections between structured SSMs and variants of attention by reformulating different approaches to computing SSMs as various matrix multiplication algorithms on structured matrices and generalizing the linear attention (Katharopoulos et al., 2020) to structured masked attention. Based on this theoretical analysis, Mamba-2 introduces an efficient SSD algorithm for SSM computation by employing block decompositions of particular semiseparable matrices and modifies the Mamba block to be tensor parallelism friendly. This framework allows Mamba-2 to leverage system optimizations developed for Transformers, including tensor parallelism and sequence parallelism, facilitating the parallel training of large models using multiple GPUs. Moreover, unlike Transformers, Mamba-2 can be efficiently trained with variable sequence lengths. Given that semiseparable matrices are extensively studied in scientific computing, incorporating these techniques presents promising opportunities for further advancing SSMs (Dao and Gu, 2024a). Besides, the adaptation of Mamba-2 to vision tasks requires further investigation.
@@ 5.1.4. Task
5.2. Causality
In visual Transformer-based architectures, the attention matrix explicitly captures dependencies between each pair of visual tokens, enabling predictions at any location to have a global receptive field. However, in Mamba-based architectures, the S6 blocks make predictions for each token based solely on the hidden state $h_{t-1}$ and the current input $x_{t}$. The $h_{t-1}$ only incorporates information from previously scanned tokens and disregards information from other tokens. Consequently, the vanilla Mamba is a causal system where predictions rely solely on the current and previous inputs. Although the causality is naturally suitable for autoregressive models in sequential vision tasks (Ren et al., 2024a), it contradicts the nature of spatial images, resulting in a fundamental mismatch between Mamba-based architectures and common visual tasks. To address this issue, existing works have proposed to integrate multiple scanning techniques to ensure a global receptive field. While these strategies partially mitigate the undesirable causality of Mamba in vision tasks, they still encounter challenges or introduce new drawbacks that require resolution.
Firstly, while the utilization of multiple scanning techniques helps establish a global receptive field, this approach does not effectively address the underlying causality issue. Tokens at different locations are still treated inconsistently due to their varying positions within the sequences, inevitably leading to a prioritization of more recently scanned tokens. Consequently, this strategy struggles to preserve spatial relationships and inadvertently introduces undesired bias for vision tasks as a result of the scanning order. In contrast, the self-attention mechanism in Transformers consistently considers all tokens to capture dependencies, showcasing its superiority to Mamba.
Secondly, the incorporation of multiple scanning strategies leads to redundancy and sub-optimal performance. To enhance Mamba’s perception at different locations, numerous approaches incorporate different scanning techniques including different scanning directions, axes, continuity, and sampling. While incorporating more scanning strategies may increase the information captured in specific locations, it introduces a significant amount of redundant information within these scanning sequences. This redundancy can impact model efficiency and hinder effective knowledge extraction for vision tasks.
@@ 5.3. In-context Learning
Currently, Transformer models are the primary large models possessing ICL capabilities. Mamba’s efficiency and long sequence modeling ability make it a promising alternative to Transformers for ICL. Furthermore, the inherent causality of the Mamba model eliminates the need for position embeddings, which are typically a constraint in extending the ICL capabilities of Transformers. In (Park et al., 2024b), the authors investigated the ICL capabilities of Mamba compared to Transformers, focusing on regression ICL tasks and small-scale models. The results demonstrated that Mamba is capable of ICL and performs comparably to Transformers in standard regression ICL tasks. Additionally, Mamba outperforms Transformers in tasks such as sparse parity learning. However, Mamba struggles with tasks involving non-standard retrieval functionality, which Transformers can handle with ease. This limitation may stem from the memory-intensive nature of these tasks, as Mamba prioritizes efficiency over contextual expressivity by utilizing a shorter latent size (Jelassi et al., 2024). Mamba-2 demonstrates notable improvements over Mamba-1 in some of these tasks, though further investigation is needed to fully understand its performance (Dao and Gu, 2024a). Research detailed in (Grazzi et al., 2024) further demonstrated Mamba’s capabilities on more complex NLP in-context tasks and showed that Mamba’s ICL efficacy scales positively with the number of in-context examples.
@@
5.4. Trustworthiness
5.4.1. Interpretability
Some studies have provided experimental evidence to elucidate the mechanisms underlying the Mamba model in NLP, focusing on its in-context learning capabilities (Grazzi et al., 2024; Park et al., 2024b), factual recall capabilities (Sharma et al., 2024), and next-token prediction capabilities (Bachmann and Nagarajan, 2024). Additionally, other works (Cirone et al., 2024) have laid theoretical foundations for Mamba’s applications in NLP. Despite these advancements, explaining why Mamba performs effectively on visual tasks remains challenging (Jafari et al., 2024; Ali et al., 2024). In (Ali et al., 2024), the authors reinterpret Mamba layers as self-attention mechanisms, thereby shedding light on its functionality both empirically and theoretically. However, the distinct learning characteristics of visual Mamba and its parallels with other foundational architectures, such as RNNs, CNNs, and ViTs, still demand deeper interpretation.
5.4.2. Generalization
According to findings in (Long et al., 2024; Guo et al., 2024b), the hidden states in Mamba are likely to accumulate or even amplify domain-specific information, which can adversely affect its generalization performance. Moreover, the 1D scanning strategies inherent to the model may inadvertently capture domain-specific biases, and current scanning techniques often fail to address the need for domain-agnostic information processing. Addressing these challenges to better harness Mamba’s strengths while improving its generalization remains a critical area for future research.
5.4.3. Safety
Research presented in (Du et al., 2024a) demonstrated VMamba’s (Liu et al., 2024e) strengths in adversarial resilience and general robustness. However, it also identified limitations in scalability when dealing with these tasks. The study includes white-box attacks on VMamba to examine the behavior of its novel components under adversarial conditions. The findings indicated that while parameter $\Delta$ exhibits robustness, parameters $\mathbf{B}$ and $\mathbf{C}$ are susceptible to attacks. This differential vulnerability among parameters contributes to VMamba’s scalability challenges in maintaining robustness. Furthermore, the results revealed that VMamba exhibits particular sensitivity to interruptions in the continuity of its scanning trajectory and the integrity of spatial information. Enhancing the safety of visual Mamba remains an unresolved challenge.
6. Conclusion